Your sole proprietor startup company is now 3 years old and you are looking for an equity investor so you can expand production. You estimate that you need $3mm in capital, and that will result in $1.5mm in cash flow next year and sustainable growth of 12%.

What ownership percentage would you need to give to a VC if they expect a 45% return on their equity?

How would that change if the growth rate went up to 15%? Down to 10%?
Answer #1

Capital Required = $ 3 million

Sustainable Growth Rate = g = 12 %, Required Return on Equity = Required Rate of Return = 45 %

Expected Perpetual Cash Flows = $ 1.5 million

Intrinsic Firm Value = 1.5 / (0.45 - 0.12) = $ 4545454.545

Ownership % to be offered = 3000000 / 4545454.545 = 0.66 or 66 %

If Growth rate rises to 15%, then Intrinsic Firm Value = 1500000 / (0.45-0.15) = $ 5000000

Ownership % to be offered = 3000000 / 5000000 = 0.6 or 60%

If Growth rate falls to 10%, then Intrinsic Firm Value = 1500000 / (0.45-0.1) = $ 4285714.286

Ownership % to be offered = (3000000 / 4285714.286) = 0.7 or 70%